If 7x + 4 = 25, then the value of x is __________. 
 If 7x + 4 = 25, then the value of x is __________. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

If 7x + 4 = 25, then the value of x is 3.
Consider the equation
7x + 4 = 25

Transposing 4 from left hand side to right hand side it becomes -4,
7x = 25–4
7x = 21
x = 21/7
x= 3

Thanks